two coherent source of light having intensity ratio 16:1 produce an interference fringe pattern
1
Expert's answer
2020-04-15T10:10:11-0400

"I_{max}=I_1+I_2+2\sqrt{I_1}\sqrt{I_2}\cdot\cos0°=(\sqrt{I_1}+\sqrt{I_2})^2"


"I_{min}=I_1+I_2+2\sqrt{I_1}\sqrt{I_2}\cdot\cos180°=(\sqrt{I_1}-\sqrt{I_2})^2"


"\frac{I_{max}}{I_{min}}=\frac{(\sqrt{I_1}+\sqrt{I_2})^2}{(\sqrt{I_1}-\sqrt{I_2})^2}=\frac{(\sqrt{16I}+\sqrt{I})^2}{(\sqrt{16I}-\sqrt{I})^2}=\frac{25}{9}"